OpenStudy (anonymous):

A compound experiment consists of drawing at random twice without replacing from a jar with 3 red, 3 white, and 2 black marbles. What is the probability of getting first a red and then a black marble?

OpenStudy (mertsj):

Probability of red = 3/8 Probability of black= 2/7 Probability of red then black = 3/8 times 2/7=3/28

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw